# JumpStarter: Getting Started on Personal Goals with AI-Powered Context Curation

### v2

@@ 3.2. Data Representation
In our system, each context is represented as a key-value pair, where the key denotes the context’s name, and the value is its content. For instance, the pair “Location preference: Midwest of US” indicates the user’s preference for a specific location - Midwest of US.
There are two types of context in our system: global context and local context. Global context refers to context that is universally applied by default throughout the system. Specifically, it consists of all context elicited from the questions the system asks at the very beginning after the user enter their goal. The local context includes all saved working solution drafts and the context added by the user throughout their interaction with the system.
@@ 3.3. Feature Implementation
In this section, we describe the implementation details of key features of our system. We introduce each feature’s implementation in the order they appeared in Section 3.1. Note that the user’s goal and the global context are integrated into the system prompt, which we use to implement all of our key features.
@@ 3.3.4. Working Solution Drafting
JumpStarter starts by selecting the relevant context for the given task. Specifically, given the title and detailed descriptions of the task, we prompt GPT-4 to select the most relevant context from the current local context collection. The selected context keys are presented as a checklist in a pop-up window in the system. Users are allowed to de-select any context key if necessary. Users can also add any context they deem useful from a drop-down list. Note that since the local context collection is empty for John when he drafts his first subtask, no local context can be selected, and thus, no pop-up window appears in his story in Section  3.1. We adopt the CoT prompting method to achieve this feature. See the prompt we used in Appendix A.3.1.
@@
As mentioned in Section  3.1, users can click on the “Add context and regenerate” button if they attempt to iterate on a working solution draft but are not sure how. The system will then perform context elicitation for working solution iteration for the user. Similar to context elicitation for goal (see Section 3.3.1), the system will provide questions to elicit more context from users, hopefully enabling them to get a better draft. To achieve this, we prompt GPT-4 in the CoT manner. See the prompt we used in Appendix A.2. With the newly elicited context, the system then triggers the “working solution generation” feature to generate an iterated draft for users.
@@ 4. Technical Evaluation
JumpStarter helps people create a hierarchical action plan and draft personalized working solutions by detecting subtasks and curating context. In this section, we describe how we evaluated the accuracy of subtask detection and the impact of context curation on generated plans and working solutions.
@@ 4.1.1. Experiment settings
Subtask detection involves determining whether a task needs further decomposition to become actionable. With vanilla zero-shot and few-shot prompting as our baselines, we experimented with three methods for detecting subtasks with GPT-4 (see Appendix B for detailed prompts). We first experimented with few-shot prompting in the Chain-of-Thought (“CoT”) style (Wei et al., 2022). We then integrated the tree levels of the task nodes into the prompt, hypothesizing that nodes at shallower levels (e.g., the first level) would be less likely to be actionable than those at deeper levels (“Tree”). We also explored incorporating the initial solution draft into the prompt to encourage the system to suggest task decomposition depending on the quality of the draft (“Draft”). We evaluated the accuracy of these methods by constructing a test suite of 20 examples from real and diverse use cases, which were labeled by expert users. In all experiments, the temperature was set to $1$, the max tokens were set to $2048$, and top-P was set to $1$. We report the average across five runs for each experimental setting.
@@ 4.2.1. Participants and Procedure
During the study, we first introduced the expert participants to the concept of action plans and working solutions with examples. We instructed the participants that they would be using three versions of the same system for their respective goal, from start to finish. The system versions represented the three experimental conditions, which were presented in shuffled order among participants to counterbalance the learning effect (see Table 2). We demonstrated how to use each version before each participant used it. We asked the participants to generate the subtasks and working solutions exactly once. They then rated the quality on a seven-point Likert scale while providing a verbal explanation for their choice.
@@ 4.2.2. Results and Findings
In addition to context selection, context elicitation improves the quality of both subtasks and working solutions, outperforming both context dumping and context filtering conditions. Participants reported that the elicitation questions posed at the beginning “provided the right plan to start with.” (E4) For instance, E2 uploaded the fellowship requirements document as initially suggested by the system, later rating the generated subtasks a perfect 7/7. “It captured the requirements quite accurately, unlike what it gave me just now [with context filtering],” E2 noted. “The subtasks were precise, fitting the unique aspects of the fellowship I am applying to, which requires only one recommendation letter, though typically more are needed.” Similarly, elicitation questions about which state to obtain a driver’s license in (for the “get a driver license” goal) and how many people are in the team (for the “organize a team event” goal) both resulted in subtasks that were better tailored to participants’ individual situations. In addition, preference elicitation questions such as “What type of vehicle do you intend to drive?” and “When do you prefer to hold the event? Weekday or weekend? Noon or night?” prompted participants to provide answers as personalized context, ultimately resulting in more tailored working solutions that they rated highly.
@@ 5. User Study
During the study, each participant was asked to explore one personal goal they would like to achieve in the near future using each system. After finishing the exploration with each system, participants were asked to evaluate the system through questionnaires on task load, outcome satisfaction, and confidence level in taking the next steps on their goal. We also conducted semi-structured interviews to understand their experiences with each system. At the end of each study session, we conducted another interview to get feedback about their overall experience and system preferences.
@@ 5.3.4. Confidence level in taking the next steps on the goal
Participants reported that ChatGPT helped validate their thoughts with commonsense knowledge and sometimes provided surprising or useful tips. As P7 said, “Sitting down and planning things out itself is very helpful. I used ChatGPT as a cross-reference, checking to make sure I’m on the right track - thinking similarly to other people. And sometimes answers to low-level tasks covered things I did not really know, which is good.” However, they reported that they still always felt they might miss something important while using ChatGPT, whereas with JumpStarter, they feel more secure (P3, P4, P9), as P4 commented, “I love that I can break things down further if I want, so I don’t feel like I miss anything.”
In addition, JumpStarter can provide very personalized and actionable next steps that greatly increase users’ confidence in taking action on their projects. For example, P1, with the goal of starting a side job, shared that “The schedule JumpStarter helped me generate is very personalized, and I can directly use it to take real action - before, I felt worried about launching this idea as I had very limited time, now I feel like I can really start doing it.” P10 also liked that JumpStarter provided them with a specific and personalized itinerary for organizing the family reunion - “I like that it summarizes everything I saved in the previous tasks - I can use it in the real world.”
@@ 5.3.5. Tool preference
8 out of 10 participants reported they prefer to use our tool in the future, compared to ChatGPT. The main reasons given include that JumpStarter can provide more customized responses with less cognitive load. The users feel that they do not have to think hard about what information to provide (P1), are guided by the system (P2), can more easily consume the information (P3) or track the plan (P6), and can get their personal details efficiently organized, framed, and utilized (P8, P9, P10).
@@ 6. Discussion
6.1. AI-powered Context Curation in Complex Problem Solving
In JumpStarter, we studied how context curation aids people in exploring complex goals. Our studies showed that context curation is essential for devising plans and working solutions of high quality. AI-driven context curation significantly reduces users’ mental burden and enhances their exploration efficiency. Context curation in JumpStarter, besides the necessary context saving, also incorporates two key components: context elicitation and context selection.
@@ 6.1. AI-powered Context Curation in Complex Problem Solving
Users greatly appreciated the elicitation questions posted by the system. Compared to the “pull” mode in ChatGPT, where users had to identify all requirements and constraints themselves to formulate a prompt, they tend to prefer the ”push” mode in JumpStarter. In this mode, they can review questions posted by JumpStarter and select which ones to answer. This preference aligns with findings in cognitive psychology literature, suggesting that recalling information is more cognitively demanding than recognizing it (Craik and Lockhart, 1972). Additionally, when users invest more effort into creating the prompt themselves, as observed in ChatGPT, they are more likely to be critical of the responses they receive. This phenomenon exemplifies the expectation confirmation theory (Oliver, 1980), which suggests that dissatisfaction arises when results do not meet the standards anticipated based on the effort invested.
@@
Users also appreciated the system assisting them in selecting context from the pool and transparently showing it before drafting a solution, giving them a good sense of control. However, it is worth noting that a few participants mentioned their desire for the system to remember and use all the context they have provided if possible. Ideally, tools like ChatGPT would maintain and use a memory of all user-provided context throughout the chat session (Ray, 2023). Our observations of users’ interactions with ChatGPT, however, often revealed that it did not effectively use the context provided to deliver more personalized responses, or even worse, it sometimes provided conflicting responses to the user. This presents an interesting challenge in updating context accordingly—as the user progresses in their problem-solving journey, their answers for some context, such as the amount of time they want to allocate each week for LSAT preparation, might change as they realize the extent of content they need to cover. In such cases, it is crucial for the system to update the user preferences in the context pool, which we aim to explore in the future.
@@
Last but not least, a major limitation of our system is that all context curation must occur within the system. Users are still required to manually upload files or input information. In line with earlier explorations (Kaptelinin, 2003; Kersten and Murphy, 2006), it would be more ideal if the system could perform more “in-situ” context capturing, integrating content from users’ everyday tools like email inboxes and desktop files. An add-on that integrates into users’ existing workflows would be ideal, as suggested by the theory of invisible technology (Bergman and Norman, 2000). One can also imagine that such an add-on could directly place the email drafts it suggests into the email application as editable drafts, ready to be refined and sent.
@@ 7. Conclusion
In this work, we present JumpStarter, a system designed to support people in getting started with their personal goals through AI-powered context curation. Our technical evaluation indicates that context curation plays a crucial role in generating plans and working solutions of higher quality. Our user study reveals that, compared to ChatGPT, JumpStarter enables users to explore their goals more efficiently and with less mental effort. We discuss the design implications of incorporating AI-powered context curation into complex problem-solving. We also provide insights into how we might blend structured and conversational approaches to better support people’s interaction with LLMs.

### v3

@@ 3.2. Data Representation
In our system, each context is represented as a key-value pair, where the key denotes the context’s name, and the value is its content. For instance, the pair “Location preference: Midwest of US” indicates the user’s preference for a specific location—Midwest of US.
There are two types of context in our system: global context and local context. Global context refers to context that is universally applied by default throughout the system. Specifically, it consists of all context elicited from the questions the system asks at the very beginning after the user enters their goal. The local context includes all saved working solution drafts and the context added by the user throughout their interaction with the system.
@@ 3.3. Feature Implementation
In this section, we provide implementation details for the key features of our system, as introduced in Section 3.1. Note that the user’s goal and the global context are integrated into the system prompt, which we use to implement all of our key features.
@@ 3.3.4. Working Solution Drafting
JumpStarter starts by selecting the relevant context for the given task. Specifically, given the title and detailed descriptions of the task, we prompt GPT-4 to select the most relevant context from the current local context collection. The selected context keys are presented as a checklist in a pop-up window in the system. Users are allowed to de-select any context key if necessary. Users can also add any context they deem useful from a drop-down list. We adopt the CoT prompting method to achieve this feature. See the prompt we used in Appendix A.3.1.
@@
As mentioned in Section  3.1, users can click on the “Add context and regenerate” button if they attempt to iterate on a working solution draft but are not sure how. The system will then perform context elicitation for working solution iteration for the user. Similar to context elicitation for goal (see Section 3.3.1), the system provides questions to elicit more context from users, hopefully enabling them to get a better draft. To achieve this, we prompt GPT-4 in the CoT manner. See the prompt we used in Appendix A.2. With the newly elicited context, the system then triggers the “working solution generation” feature to generate an iterated draft for users.
@@ 4. Technical Evaluation
JumpStarter helps people create a hierarchical action plan and draft working solutions by detecting subtasks and curating personal context. In this section, we describe how we evaluated the accuracy of subtask detection and the impact of context curation on generated plans and working solutions.
@@ 4.1.1. Experiment settings
Subtask detection involves determining whether a task needs further decomposition to become actionable. With vanilla zero-shot and few-shot prompting as our baselines, we experimented with three methods for detecting subtasks with GPT-4 (see Appendix B for detailed prompts). We first experimented with few-shot prompting in the Chain-of-Thought (“CoT”) style (Wei et al., 2022). We then integrated the tree levels of the task nodes into the prompt, hypothesizing that nodes at shallower levels (e.g., the first level) would be less likely to be actionable than those at deeper levels (“Tree”). We also explored incorporating the initial solution draft into the prompt to encourage the system to suggest task decomposition depending on the quality of the draft (“Draft”). We evaluated the accuracy of these methods on an expert-labeled test suite. In all experiments, the temperature was set to $1$, the max tokens were set to $2048$, and top-P was set to $1$. We report the average across five runs for each experimental setting.
4.1.2. Test suite construction
Drawing inspiration from (Little, 1983), we created a test suite comprising four real-world task scenarios, each representing a distinct aspect of everyday life: 1) Applying for a PhD program (Academic), 2) Obtaining a driver’s license (Practical), 3) Finding a surfing camp (Recreational), and 4) Arranging a trip abroad (Travel). For each scenario, we created five distinct test cases, resulting in a total of 20 test cases. An example test case is shown below:
Compile a List of Potential Universities: start by identifying the universities that offer PhD programs in Natural Language Processing (NLP). Research and compile a comprehensive list of these universities.
We recruited four expert participants—one for each task scenario—via a university mailing list (average age=26.3, two female, two male). These participants reported having completed the tasks in the past six months and were compensated \$10 for their participation. For each test case, the experts were asked to determine if the current task should be further decomposed to make it actionable. For the example test case shown above, its label is No as it is actionable and does not require further task decomposition.
@@ 4.2.1. Participants and Procedure
During the study, we first introduced the expert participants to the concept of action plans and working solutions with examples. We instructed the participants that they would be using three versions of the same system for their respective goal, from start to finish. The system versions represented the three experimental conditions, which were presented in shuffled order among participants to counterbalance the learning effect (see Table 2). We demonstrated how to use each version before each participant used it. We asked the participants to generate the subtasks and working solutions exactly once. They then rated the perceived quality of the subtasks and working solutions on a seven-point Likert scale, providing a verbal explanation for their choices.
@@ 4.2.2. Results and Findings
In addition to context selection, context elicitation improves the quality of both subtasks and working solutions, outperforming both context dumping and context filtering conditions. Participants reported that the elicitation questions posed at the beginning “provided the right plan to start with.” (E4) For instance, E2 uploaded the fellowship requirements document as initially suggested by the system, later rating the generated subtasks a perfect 7/7. “It captured the requirements quite accurately…” E2 noted. “The subtasks were precise, fitting the unique aspects of the fellowship I am applying to, which requires only one recommendation letter, though typically more are needed.” Similarly, elicitation questions about which state to obtain a driver’s license in (for the “get a driver license” goal) and how many people are in the team (for the “organize a team event” goal) both resulted in subtasks that were better tailored to participants’ individual situations. In addition, preference elicitation questions such as “What type of vehicle do you intend to drive?” and “When do you prefer to hold the event? Weekday or weekend? Noon or night?” prompted participants to provide answers as personal context, ultimately resulting in more tailored working solutions that they rated highly.
@@ 5. User Study
During the study, each participant was asked to explore one personal goal they would like to achieve in the near future using each system. After finishing the exploration with each system, participants were asked to evaluate the system through questionnaires on task load, outcome satisfaction, and confidence level in taking the next steps on their goal. We also conducted semi-structured interviews to understand their experiences with each system. At the end of each study, we conducted another interview to get feedback about their overall experience and system preferences.
@@ 5.3.4. Confidence level in taking the next steps on the goal
Participants reported that ChatGPT helped validate their thoughts with commonsense knowledge and sometimes provided surprising or useful tips. As P7 said, “Sitting down and planning things out itself is very helpful. I used ChatGPT as a cross-reference, checking to make sure I’m on the right track—thinking similarly to other people. And sometimes answers to low-level tasks covered things I did not really know, which is good.” However, they reported that they still always felt they might miss something important while using ChatGPT, whereas with JumpStarter, they feel more secure (P3, P4, P9), as P4 commented, “I love that I can break things down further if I want, so I don’t feel like I miss anything.”
In addition, JumpStarter can provide very personalized and actionable next steps that greatly increase users’ confidence in taking action on their projects. For example, P1, with the goal of starting a side job, shared that “The schedule JumpStarter helped me generate is very personalized, and I can directly use it to take real action—before, I felt worried about launching this idea as I had very limited time, now I feel like I can really start doing it.” P10 also liked that JumpStarter provided them with a specific and personalized itinerary for organizing the family reunion—“I like that it summarizes everything I saved in the previous tasks—I can use it in the real world.”
@@ 5.3.5. Tool preference
8 out of 10 participants reported they prefer to use JumpStarter in the future, compared to ChatGPT. The main reasons given include that JumpStarter can provide more customized responses with less cognitive load. The users feel that they do not have to think hard about what information to provide (P1), are guided by the system (P2), can more easily consume the information (P3) or track the plan (P6), and can get their personal details efficiently organized, framed, and utilized (P8, P9, P10).
@@ 6. Discussion
6.1. Adpative Personal Context Curation in Complex Problem Solving
JumpStarter’s use of adaptive personal context curation in helping users explore complex goals provides a design paradigm for future interfaces. Systems that effectively respond to personal context can better collaborate with users in achieving their goals, as demonstrated in our user studies. In JumpStarter, context curation involves not only saving necessary information but also incorporates two key components: context elicitation and context selection. We found that soliciting and integrating relevant context is essential for devising personalized plans and solutions of high quality.
@@ 6.1. Adpative Personal Context Curation in Complex Problem Solving
Users greatly appreciated the elicitation questions posted by the system. Compared to the “pull” mode in ChatGPT, where users had to identify the requirements and constraints themselves, they tend to prefer the ”push” mode in JumpStarter, where they can review the posted questions and select which ones to answer. This preference aligns with cognitive psychology literature, suggesting that recalling information is more cognitively demanding than recognizing it (Craik and Lockhart, 1972). Additionally, when users invest more effort into prompting, as observed in ChatGPT, they are more likely to be critical of the responses they receive. This exemplifies the expectation confirmation theory (Oliver, 1980), which suggests that dissatisfaction arises when results do not meet the standards anticipated based on the effort invested.
@@
Users also appreciated the system’s assistance in selecting and previewing context before drafting a solution, giving them a good sense of transparency and control. However, a few participants expressed a desire for the system to remember and use all the context they had provided. Tools like ChatGPT are designed to maintain and utilize a memory of all user-provided context throughout the chat session (Ray, 2023). However, our observations reveal that ChatGPT did not effectively use the provided context to deliver more personalized responses, and sometimes even provided conflicting responses to the user. This presents an interesting challenge  (Wang et al., 2024)—as users progress in their problem-solving journey, their answers to some context, such as the amount of time they want to allocate each week for LSAT preparation, might change as they realize the extent of content they need to cover. In such cases, it is crucial for the system to update the user preferences in the context pool. We aim to explore ways to refine context selection and integration from an existing context pool in the future.
@@
Adaptive personal context curation advances the vision of flexible and personalized systems that actively respond to user intent. Currently, in JumpStarter, all context curation occurs within the system—users must manually upload files or input information rather than integrating content from existing tools like email inboxes and desktop files. If more fully integrated into users’ daily workflows, JumpStarter could solicit and integrate more relevant context, providing even more actionable and useful support to users. In line with earlier explorations (Kaptelinin, 2003; Kersten and Murphy, 2006), it would be ideal if the system could perform “in-situ” context capturing, such as an add-on that integrates into users’ existing workflows, as suggested by the theory of invisible technology (Bergman and Norman, 2000). Such an add-on could directly place the email drafts it suggests into the email application as editable drafts, ready to be refined and sent. Ultimately, JumpStarter’s use of adaptive personal context curation provides a design paradigm for future systems, where human insight and computational assistance can be more effectively blended together to achieve goals.
@@ 6.3. Limitations and Future Work
JumpStarter supports users from planning to initiating their projects by creating actionable prototypes. This helps users grasp the full scope of their tasks and make tangible progress. While JumpStarter does not perform tasks like sending emails, it prepares and organizes essential information for action. This initial drafting stage is crucial for overcoming the planning fallacy, where users often underestimate the complexity and time required for tasks, leading to execution difficulties (Ahmetoglu et al., 2021). Moving forward, our focus will be on assisting users throughout their project’s lifespan. Future research will explore strategies for maintaining momentum and adapting to changes during real-world execution.
@@ 7. Conclusion
In this work, we present JumpStarter, a system designed to support people in getting started with their personal goals through adaptive personal context curation. Our technical evaluation indicates that context curation plays a crucial role in generating plans and working solutions of higher quality. Our user study reveals that, compared to ChatGPT, JumpStarter enables users to explore their goals more efficiently and with less mental effort. We discuss the design implications of incorporating adaptive personal context curation into complex problem-solving. We also provide insights into how we might blend structured and conversational approaches to better support people’s interaction with LLMs.